OpenStudy (anonymous):

A large pond is stocked with fish. The fish population P is modeled by the formula P = 2t + 10(sqrt of t) + 120, where t is the number of days since the fish were first introduced into the pond. How many days will it take for the fish population to reach 500? (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

You want the population to be 500. We can insert this into the equation: \[500 = 2t + 10\sqrt{t}+120\] do you know what to do now?

OpenStudy (anonymous):

no where do i go from here

OpenStudy (anonymous):

You can first subtract 120 from both sides... \[500-120 = 2t + 10\sqrt{t} + 120 - 120\] Simplify... \[380 = 2t + 10\sqrt{t}\]

OpenStudy (anonymous):

then whats it come out too

OpenStudy (anonymous):

120?

OpenStudy (anonymous):

peddani?